OpenStudy (anonymous):

ln x-ln(x-2)=1 find the value of x

OpenStudy (anonymous):

ln (x) - ln(x-2) = 1 ln(x/x-2) = 1 e(x-2) = x (e-1)x = 2e x = 3.16

OpenStudy (anonymous):

how do i plugged that in the calulator